A stone is dropped from the top of a two-story building said to be 855 ft. from the ground. (a) How long does it take the stone to reach this height? (b) Compute for the rate of the stone’s motion upon reaching the ground. Show your solution.
Answer by Alan3354(69443)

A stone is dropped from the top of a two-story building said to be 855 ft. from the ground. (a) How long does it take the stone to reach this height? (b) Compute for the rate of the stone’s motion upon reaching the ground. Show your solution.
-------------------
That's a tall building for only 2 stories.
---
A stone is dropped from the top of a two-story building said to be 855 ft. from the ground. (a) How long does it take the stone to reach this height?
IDK. Does it take an elevator or go up the stairs?
Maybe you mean to reach the ground?
Using 32 ft/sec/sec for gravity:
s = at^2/2
855 = 32t^2/2
t^2 = 855/16
t =~ 7.311 seconds
================================
(b) Compute for the rate of the stone’s motion upon reaching the ground.
v = a*t = 32*7.311 = ~233.9 ft/sec